In a certain town, male and female each form 50% of the population. It is known that 20% of the males and 5% of the females are unemployed. A research student studying the employment situation selects an unemployed person at random. What is the probability that the person so selected is: (a) male (b) Female

Answer by greenestamps(13216)



Given that the numbers of males and females are the same, the 20% of males that are unemployed and the 5% of females that are unemployed means there are 4 times as many males as females that are unemployed -- i.e., 4/5 of the unemployed people are male. So

(a) P(randomly selected unemployed person is male) = 4/5 = 80% = 0.80
(b) P(randomly selected unemployed person is female) = 1/5 = 20% = 0.20